# Cygnus X-3 as semi-hidden PeVatron

### v1

@@ Abstract
The high-mass X-ray binary Cygnus X-3 has been suggested for a long time to be a source of high-energy photons and neutrinos. In view of the increased sensitivity of current experiments, we examine the acceleration and interactions of high-energy cosmic rays (CRs) in this binary system, assuming that the compact object is a black hole. Using a test-particle approach in a Monte-Carlo framework, we employ as the basic CR acceleration mechanisms magnetic reconnection or 2.nd order Fermi acceleration and diffusive shock acceleration. We find that in all three scenarios CRs can be accelerated beyond PeV energies. High-energy photons and neutrinos are produced as secondaries in photo-hadronic interactions of CRs on X-ray photons and in the scattering on gas from the wind of the companion star. Normalising the predicted photon flux to the excess flux observed by LHAASO at energies above PeV in the direction of Cygnus X-3, a CR acceleration efficiency of $10^{-3}$ is sufficient to power the required CR luminosity. Our results suggest that the PeV photon flux from Cygnus X-3  could be in a bright phase significantly increased relative to the average flux of the last years.
Subject headings:
@@
1. Introduction
Cygnus X-3, a high-mass X-ray binary in the direction of the Cygnus OB association, has had an outstanding impact on the development of gamma-ray astronomy. In 1983, Samorski & Stamm found a periodically modulated flux of neutral particles with energies above $2\times 10^{15}$ eV coincident with Cygnus X-3 in archival data from the Kiel air-shower array. This apparent detection and the supporting evidence provided during the following years by various experiments operating in the northern hemisphere triggered an enormous theoretical interest in this source (Gaisser & Stanev, 1985; Berezinsky et al., 1985, 1986). More importantly, these early results on Cygnus X-3 stimulated the building of air-shower arrays designed specifically for gamma-ray astronomy. However, none of these newly built arrays, like CASA-MIA and SPASE, detected a signal from Cygnus X-3, for a review of these early results see Protheroe (1994). Moreover, theoretical arguments suggest that most of these claims, in particular those by underground experiments, were erroneous. Still, it is a tantalising option that the initial claims for a photon signal were correct. In this case, Cygnus X-3 might have been in a bright phase in the early 1980s, which ended before the start of data-taking of the following generation of experiments. Such a strong variability is consistent with the decrease by at least a factor of 100 reported in the long-term average flux between the mid-1970s and mid-1980s (Bhat et al., 1986).
@@ 2. Environment of Cyg X-3
The determination of radial velocity curves for the Cygnus X-3 binary system is complicated by strong optical extinction and the difficult separation of wind features from those arising in the photosphere of the companion. Therefore, the individual masses of the components, $M_{\rm CO}$ of the compact object and $M_{\rm WR}$ of the companion Wolf-Rayet (WR) star, have remained uncertain up-to-date. Published results span a wide range from a BH with mass close to or above $20\,M_{\odot}$ (Schmutz et al., 1996; Hjalmarsdotter et al., 2008) down to upper limits of $3.6\,M_{\odot}$ (Stark & Saia, 2003). In contrast, the mass ratio of the binary stars can be constrained more precisely as $M_{\rm CS}/M_{\rm WR}=0.24\pm 0.06$ combining radial velocity curves derived from FeXXVI emission lines with infrared HeI absorption lines (Hanson et al., 2000). Luckily, several parameters important for the determination of the spectra of high-energy particles have a rather weak dependence on the masses of the binary system. For instance, the orbital separation $a$, which in turn influences the target density of stellar photons and the gas in the stellar wind, depends only as $a\propto(M_{\rm CS}+M_{\rm WR})^{1/3}$ on the total mass of the binary system. To be able to compare easier with earlier studies, we will use values close to the upper mass range discussed in the literature, setting $M_{\rm CS}=M_{\rm BH}=20M_{\odot}$ and $M_{\rm WR}=50M_{\odot}$ corresponding to the high-mass solution of  Szostek & Zdziarski (2008). We will disuses later how our results change if these masses are reduced. Thus the Schwarzschild radius of the BH is $R_{s}=6\times 10^{6}\,$cm, the orbital separation of the stars in the binary system is $a=3\times 10^{11}$ cm, while we use 9.6 kpc as distance to Cygnus X-3 (Reid & Miller-Jones, 2023).
@@
Figure 1.— Spectral density of photon backgrounds in the state $S_{1}$ (left) and $S_{2}$ (right).
@@
Since the radial extension of the disk is rather small, the total emission is close to a Planck distribution with $T(r)\simeq T(R_{0})$. For a helium-rich composition and the standard value $\eta=0.1$ for the accretion efficiency $\eta=L_{\rm Edd}/(\dot{M}_{\rm Edd}c^{2})$, the Eddington luminosity becomes $L_{\rm Edd}\simeq 2.6\times 10^{38}(M_{\rm BH}/M_{\odot})$erg/s which in turn fixes $\dot{M}=\lambda\dot{M}_{\rm Edd}$ for a chosen $\lambda$. The observed X-ray luminosity favours large value of $\lambda$ and $M_{\rm BH}$ (Veledina et al., 2024). We calculate the density of disk photons at a given point in the jet integrating the photon emissivity into the corresponding solid angle ${\rm d}\Omega$ over the accretion disk (Cerutti et al., 2011), choosing $\lambda=0.5$ and as smallest radius $R_{0}$ of the acceleration and of the emission region six Schwarzschild radii, $R_{0}=6R_{s}\simeq 3.6\times 10^{7}\,$cm, while we set $R_{\rm out}=10R_{s}$ for the outer radius of the disk.
@@
3. Theoretical framework
3.1. Acceleration scenarios
@@
$t_{\rm Fermi}^{-1}=\frac{1}{E}\frac{{\rm d}E}{{\rm d}t}_{\rm Fermi}\simeq 4\kappa(\Gamma\beta)^{2}\frac{L_{c}}{v_{A}}\simeq 10^{5}{\rm s}^{-1}.$ (6)
@@
Figure 2.— Rates for synchrotron losses $t^{-1}_{\rm syn}$, photo-hadronic $t^{-1}_{p\gamma}$ and hadronic $t^{-1}_{pp}$ interactions of proton as function of energy compared to the acceleration rate by reconnection $t^{-1}_{\rm rec}$ and 2.nd order Fermi acceleration $t^{-1}_{\rm Fermi}$ in the state $S_{1}$ (left) and by DSA $t^{-1}_{\rm sh}$ in the state $S_{2}$ (right).
3.2. Monte Carlo scheme
@@
Here, $\alpha$ is the theoretically predicted slope of the differential energy spectrum of accelerated particles in the case of no interactions and energy losses, which we choose for both acceleration mechanisms as $\alpha=2.3$, while $\xi=({\rm d}E/{\rm d}t)_{\rm acc}\Delta t/E$ is the energy fraction gained according to Eqs. (4) and (5). If the particle stays in the acceleration process, the true energy gain is calculated as the sum of the acceleration gain and the continuous energy losses, $[({\rm d}E/{\rm d}t)_{\rm acc}-({\rm d}E/{\rm d}t)_{\rm CEL}]\Delta t$, where the latter include synchrotron and, for muons, inverse Compton losses. Finally, it is decided if the particle decays or interacts.
@@
If a photo-hadronic interactions occurs, we use the modified version of the SOPHIA program (Mucke et al., 2000) developped by Kachelrieß et al. (2008) to generate secondary particles, while we employ QGSJET-IIc (Ostapchenko, 2011, 2013) for hadronic interactions. We take into account the finite decay time of all unstable particle except neutral pions and include only stable charged all particles in the acceleration process. For the decay of unstable particles, we employ SIBYLL (Fletcher et al., 1994).
@@
After the particles escaped from the acceleration region, we assume that the magnetic field strength is small enough that they move semi-ballistically. Except for photons, the interaction depth is small, $\tau\ll 1$, and multiple interactions can be neglected. Thus we add after escape a final interaction for hadrons with probability $p=1-\exp(-\tau)$ and let all unstable particles decay. In the case of photons, we calculate the pair-production probability on background photons in the direction of the observer. In addition, we add the pair-production probability on photons of the cosmic microwave background (CMB) and the extragalactic background light from Franceschini et al. (2008) to obtain the final flux observed on Earth, $F=F_{0}\exp(-\tau_{\gamma\gamma})$. We neglect the absorption on the starlight in the Milky Way, which adds only a minor correction relative to the absorption on the CMB (Vernetto & Lipari, 2016) and the effect of electromagnetic cascades what is well justified above the pair-production threshold, $E_{\gamma}\simeq(1-10)$ GeV.
@@
Note that the acceleration time is much smaller than the orbital period of the system. Thus there is an orbital modulation of the fluxes, both because the interaction depth varies with the orbital phase and, in addition for photons, because the absorption probability changes. In the energy range we are mainly interested in, $10^{14}$$10^{16}$ eV, the modulation is however only minor: The photon absorption probability is dominated by the CMB, while the interaction depths vary only by a factor a few, cf. Lammert (2025) for details. In the following, we present therefore results for the average fluxes.
4. Numerical results
@@
Figure 3.— Particle fluxes of CR protons, sum of all neutrino flavours and photon fluxes as function of energy $E$ for the case of reconnection in the state $S_{1}$ (left) and DSA in the state $S_{2}$ (right).
4.1. Proton injection
@@
We normalise the fluxes such that the CR luminosity satisfies $L_{\rm CR}=\eta_{\rm CR}\lambda L_{\rm Edd}\simeq 3.9\times 10^{38}$erg/s, i.e. we assume optimistically that 15% of the available energy is used to accelerate hadrons. In Fig. 3, we show the rescaled flux $E^{2}\Phi=E^{2}{\rm d}N/({\rm d}A{\rm d}t{\rm d}E)$ of CR protons (blue line), of the sum of all neutrino flavours (orange line) and of photons (red line) as function of energy $E$. In addition to the proton flux after interactions, the pressumed power-law $E^{-2.3}$ and the flux escaping the acceleration region are presented. The high-energy cutoffs agree with the expectations from Fig. 2, i.e. a rather soft suppression above $8\times 10^{16}$  eV in the state $S_{1}$ and a sharper suppression above $2\times 10^{16}$  eV in the state $S_{2}$. Similarly, the secondary fluxes in the state $S_{1}$ are as expected larger than in the state $S_{2}$. The red dashed line indicates the unabsorbed photon flux, while the solid line shows the photon flux taking into account pair production: At energies above $10^{14}$ eV pair production on CMB photons dominates, while at lower energies the absorption on stellar photons becomes important.
@@
In Fig. 4, we show a close-up of the photon fluxes, where we split the flux into two components, depending on their production inside or outside the acceleration zone. The outer component is mainly produced in hadronic interactions, and thus the slope of this component repeats the slope of the proton flux at an energy a factor $\simeq 20$ higher. In addition, we show in Fig. 4 the gamma-ray flux measured by LHAASO in the direction of the Cygnus X superbubble from Cao et al. (2024). Since no detailed information on the exposure is published, we can only roughly estimate the integrated flux $\Phi(>E)$ from the central region above PeV energies as $\Phi(>{\rm PeV})=N/(AT)\simeq 9\times 10^{-14}/({\rm m^{2}s})$ using as effective area $A\simeq 10^{6}$m2 from Ma et al. (2022), $T\simeq 6000$ h for 3 years of data taking, and $N=2$ events above PeV. This corresponds to a photon flux of order $E^{2}\Phi(E={\rm PeV})\simeq 10^{-11}$eV/cm2s. Thus the obtained photon flux, which is in both states on the level $E^{2}\Phi(E={\rm PeV})\simeq 10^{-9}$eV/cm2s, should be scaled down by $10^{-2}$, what would reduce the required CR acceleration efficiency down to $\eta_{\rm CR}\sim 10^{-3}$. At this level, the photon flux would be well below the lower limits from the MAGIC collaboration (Aleksić et al., 2010). The correspondingly down-scaled neutrino flux is a factor $100$ below the 90% C.L. upper limit set by the IceCube collaboration (Abbasi et al., 2022). Even for $\eta_{\rm CR}=0.1$, the number of expected muon neutrino events/year in IceCube above 10 TeV is only around one in $S_{1}$ and two in $S_{2}$.
@@
Figure 4.— Photon fluxes as function of energy $E$ for the case of reconnection in the state $S_{1}$ (left) and by DSA in the state $S_{2}$ (right).
4.2. Helium injection

### v2

@@ Abstract
Context. The high-mass X-ray binary Cygnus X-3 has been suggested for a long time to be a source of high-energy photons and neutrinos.
Aims. In view of the increased sensitivity of current experiments, we examine the acceleration and interactions of high-energy cosmic rays (CRs) in this binary system, assuming that the compact object is a black hole.
Methods. Using a test-particle approach in a Monte-Carlo framework, we employ as the basic CR acceleration mechanisms magnetic reconnection or 2.nd order Fermi acceleration and diffusive shock acceleration.
Results. We find that in all three scenarios CRs can be accelerated beyond PeV energies. High-energy photons and neutrinos are produced as secondaries in photo-hadronic interactions of CRs on X-ray photons and in the scattering on gas from the wind of the companion star. Normalising the predicted photon flux to the excess flux observed by LHAASO at energies above PeV in the direction of Cygnus X-3, a CR acceleration efficiency of $10^{-3}$ is sufficient to power the required CR luminosity. Our results suggest that the PeV photon flux from Cygnus X-3  could be in a bright phase significantly increased relative to the average flux of the last years.
Key Words.:
@@
1 Introduction
Cygnus X-3, a high-mass X-ray binary in the direction of the Cygnus OB association, has had an outstanding impact on the development of gamma-ray astronomy. In 1983, Samorski & Stamm found a periodically modulated flux of neutral particles with energies above $2\times 10^{15}$ eV coincident with Cygnus X-3 in archival data from the Kiel air-shower array. This apparent detection and the supporting evidence provided during the following years by various experiments operating in the northern hemisphere triggered an enormous theoretical interest in this source (Gaisser & Stanev, 1985; Berezinsky et al., 1985, 1986). More importantly, these early results on Cygnus X-3 stimulated the building of air-shower arrays designed specifically for gamma-ray astronomy. However, none of these newly built arrays, like CASA-MIA and SPASE, detected a signal from Cygnus X-3, for a review of these early results see Protheroe (1994). Moreover, theoretical arguments suggest that most of these claims, in particular those by underground experiments, were erroneous as discussed already by Berezinsky et al. (1986). Still, it is a tantalising option that the initial claims for a photon signal were correct. In this case, Cygnus X-3 might have been in a bright phase in the early 1980s, which ended before the start of data-taking of the following generation of experiments. Such a strong variability is consistent with the decrease by at least a factor of 100 reported in the long-term average flux between the mid-1970s and mid-1980s (Bhat et al., 1986).
@@ 2 Environment of Cyg X-3
The determination of radial velocity curves for the Cygnus X-3 binary system is complicated by strong optical extinction and the difficult separation of wind features from those arising in the photosphere of the companion. Therefore, the individual masses of the components, $M_{\rm CO}$ of the compact object and $M_{\rm WR}$ of the companion Wolf-Rayet (WR) star, have remained uncertain up-to-date. Published results span a wide range from a BH with mass close to or above $20\,M_{\odot}$ (Schmutz et al., 1996; Hjalmarsdotter et al., 2008) down to upper limits of $3.6\,M_{\odot}$ (Stark & Saia, 2003). In contrast, the mass ratio of the binary stars can be constrained more precisely as $M_{\rm CO}/M_{\rm WR}=0.24\pm 0.06$ combining radial velocity curves derived from FeXXVI emission lines with infrared HeI absorption lines (Hanson et al., 2000). Luckily, several parameters important for the determination of the spectra of high-energy particles have a rather weak dependence on the masses of the binary system. For instance, the orbital separation $a$, which in turn influences the target density of stellar photons and the gas in the stellar wind, depends only as $a\propto(M_{\rm CO}+M_{\rm WR})^{1/3}$ on the total mass of the binary system. To be able to compare easier with earlier studies, we will use values close to the upper mass range discussed in the literature, setting $M_{\rm CO}=M_{\rm BH}=20M_{\odot}$ and $M_{\rm WR}=50M_{\odot}$ corresponding to the high-mass solution of  Szostek & Zdziarski (2008). We will disuses later how our results change if these masses are reduced. Thus the Schwarzschild radius of the BH is $R_{s}=6\times 10^{6}\,$cm, the orbital separation of the stars in the binary system is $a=3\times 10^{11}$ cm, while we use 9.6 kpc as distance to Cygnus X-3 (Reid & Miller-Jones, 2023).
@@
Figure 1: Spectral density of photon backgrounds in the state $S_{1}$ (left) and $S_{2}$ (right).
@@
Since the radial extension of the disk is rather small, the total emission is close to a Planck distribution with $T(r)\simeq T(R_{0})$. For a helium-rich composition and the standard value $\eta=0.1$ for the accretion efficiency $\eta=L_{\rm Edd}/(\dot{M}_{\rm Edd}c^{2})$, the Eddington luminosity becomes $L_{\rm Edd}\simeq 2.6\times 10^{38}(M_{\rm BH}/M_{\odot})$erg/s which in turn fixes $\dot{M}=\lambda\dot{M}_{\rm Edd}$ for a chosen $\lambda$. The observed X-ray luminosity favours a large value of $\lambda$ and $M_{\rm BH}$ (Veledina et al., 2024). We calculate the density of disk photons at a given point in the jet integrating the photon emissivity into the corresponding solid angle ${\rm d}\Omega$ over the accretion disk (Cerutti et al., 2011), choosing $\lambda=0.5$ and as smallest radius $R_{0}$ of the acceleration and of the emission region six Schwarzschild radii, $R_{0}=6R_{s}\simeq 3.6\times 10^{7}\,$cm, while we set $R_{\rm out}=10R_{s}$ for the outer radius of the disk.
@@
3 Theoretical framework
3.1 Acceleration scenarios
@@
$t_{\rm Fermi}^{-1}=\frac{1}{E}\frac{{\rm d}E}{{\rm d}t}_{\rm Fermi}\simeq 4\kappa(\Gamma\beta)^{2}\frac{v_{A}}{L_{c}}\simeq 10^{5}{\rm s}^{-1}.$ (6)
@@
Figure 2: Rates for synchrotron losses $t^{-1}_{\rm syn}$, photo-hadronic $t^{-1}_{p\gamma}$ and hadronic $t^{-1}_{pp}$ interactions of proton as function of energy compared to the acceleration rate by reconnection $t^{-1}_{\rm rec}$ and 2.nd order Fermi acceleration $t^{-1}_{\rm Fermi}$ in the state $S_{1}$ (left) and by DSA $t^{-1}_{\rm sh}$ in the state $S_{2}$ (right).
3.2 Monte Carlo scheme
@@
Here, $\alpha$ is the theoretically predicted slope of the differential energy spectrum of accelerated particles in the case of no interactions and energy losses, which we choose for both acceleration mechanisms as $\alpha=2.3$, while $\xi=({\rm d}E/{\rm d}t)_{\rm acc}\Delta t/E$ is the energy fraction gained according to Eqs. (4) and (5). If the particle stays in the acceleration process, the true energy gain is calculated as the sum of the acceleration gain and the continuous energy losses, $[({\rm d}E/{\rm d}t)_{\rm acc}-({\rm d}E/{\rm d}t)_{\rm CEL}]\Delta t$, where the latter include synchrotron and, for muons, inverse Compton losses. Finally, it is decided if the particle decays or interacts: First, the probability that something happens is compared to a random number $r$ uniformly distributed in [0:1[[0:1[[ 0 : 1 [, then a second random number $r$ is compared to the branching ratios for the relevant decay and interaction channels.
@@
If a photo-hadronic interactions occurs, we use the modified version of the SOPHIA program (Mucke et al., 2000) developped by Kachelrieß et al. (2008) to generate secondary particles, while we employ QGSJET-IIc (Ostapchenko, 2011, 2013) for hadronic interactions. We take into account the finite decay time of all unstable particles except neutral pions and include only stable charged all particles in the acceleration process. For the decay of unstable particles, we employ SIBYLL (Fletcher et al., 1994).
@@
After the particles escaped from the acceleration region, we assume that the magnetic field strength is small enough that they move ballistically. Except for photons, the interaction depth is small, $\tau\ll 1$, and multiple interactions can be neglected. Thus we add after escape a final interaction for hadrons with probability $p=1-\exp(-\tau)$ and let all unstable particles decay. In the case of photons, we calculate the pair-production probability on background photons in the direction of the observer. In addition, we add the pair-production probability on photons of the cosmic microwave background (CMB) and the extragalactic background light from Franceschini et al. (2008) to obtain the final flux observed on Earth, $F=F_{0}\exp(-\tau_{\gamma\gamma})$. We neglect the absorption on the starlight in the Milky Way, which adds only a minor correction relative to the absorption on the CMB (Vernetto & Lipari, 2016) and the effect of electromagnetic cascades what is well justified above the pair-production threshold, $E_{\gamma}\simeq(1-10)$ GeV.
@@
Note that the acceleration time is much smaller than the orbital period of the system. Thus there is an orbital modulation of the fluxes, both because the interaction depth varies with the orbital phase and, in addition for photons, because the absorption probability changes. In the energy range we are mainly interested in, $(10^{14}$$10^{16})$ eV, the modulation is however only minor: The photon absorption probability is dominated by the CMB, while the interaction depths vary only by a factor a few, cf. Lammert (2025) for details. In the following, we present therefore results for the average fluxes.
4 Numerical results
@@
Figure 3: Particle fluxes of CR protons, sum of all neutrino flavours and photon fluxes as function of energy $E$ for the case of reconnection in the state $S_{1}$ (left) and DSA in the state $S_{2}$ (right).
4.1 Proton injection
@@
We normalise the fluxes such that the CR luminosity satisfies $L_{\rm CR}=\eta_{\rm CR}\lambda L_{\rm Edd}\simeq 3.9\times 10^{38}$erg/s, i.e. we assume optimistically that 15% of the available energy is used to accelerate hadrons. In Fig. 3, we show the rescaled flux $E^{2}\Phi=E^{2}{\rm d}N/({\rm d}A{\rm d}t{\rm d}E)$ of CR protons (blue line), of the sum of all neutrino flavours (orange line) and of photons (red line) as function of energy $E$. In addition to the proton flux after interactions, the pressumed power-law $E^{-2.3}$ and the flux escaping the acceleration region are presented. The high-energy cutoffs agree with the expectations from Fig. 2, i.e. a rather soft suppression above $8\times 10^{16}$  eV in the state $S_{1}$ and a sharper suppression above $2\times 10^{16}$  eV in the state $S_{2}$. Similarly, the secondary fluxes in the state $S_{1}$ are as expected larger than in the state $S_{2}$. The red dotted line indicates the unabsorbed photon flux, while the dot-dashed line shows the photon flux taking into account pair production: At energies above $10^{14}$ eV pair production on CMB photons dominates, while at lower energies the absorption on stellar photons becomes important.
@@
In Fig. 4, we show a close-up of the photon fluxes, where we split the flux into two components, depending on their production inside or outside the acceleration zone. The outer component is mainly produced in hadronic interactions, and thus the slope of this component repeats the slope of the proton flux at an energy a factor $\simeq 20$ higher. In addition, we show in Fig. 4 the gamma-ray flux measured by LHAASO in the direction of the Cygnus X superbubble from Cao et al. (2024). Note that Cygnus X-3 contributes only a fraction, mostly at the highest energies, of the total flux from the Cygnus superbubble. Since no detailed information on the exposure is published, we can only roughly estimate the integrated flux $\Phi(>E)$ from the central region above PeV energies as $\Phi(>{\rm PeV})=N/(AT)\simeq 9\times 10^{-14}/({\rm m^{2}s})$ using as effective area $A\simeq 10^{6}$m2 from Ma et al. (2022), $T\simeq 6000$ h for 3 years of data taking, and $N=2$ events above PeV. This corresponds to a photon flux of order $E^{2}\Phi(E={\rm PeV})\simeq 10^{-11}$eV/cm2s. Thus the obtained photon flux, which is in both states on the level $E^{2}\Phi(E={\rm PeV})\simeq 10^{-9}$eV/cm2s, should be scaled down by $10^{-2}$, what would reduce the required CR acceleration efficiency down to $\eta_{\rm CR}\sim 10^{-3}$. At this level, the photon flux would be well below the lower limits from the MAGIC collaboration (Aleksić et al., 2010). The correspondingly down-scaled neutrino flux is a factor $100$ below the 90% C.L. upper limit set by the IceCube collaboration (Abbasi et al., 2022). Even for $\eta_{\rm CR}=0.1$, the number of expected muon neutrino events/year in IceCube above 10 TeV is only around one in $S_{1}$ and two in $S_{2}$.
@@
Figure 4: Photon fluxes as function of energy $E$ for the case of reconnection in the state $S_{1}$ (left) and by DSA in the state $S_{2}$ (right) compared to the gamma-ray flux measured by LHAASO (Cao et al., 2024) in the direction of the Cygnus X superbubble.
4.2 Helium injection